OpenStudy (anonymous):

solve g^x+2 = (1/27) ^x=12

OpenStudy (anonymous):

Okay, so you first take (1/27)^x=12 Taking log on both sides, you have xlog(1/27)=log(12) That ix, -3xlog(3)=log(12) *I trust you're aware how) Which gives me, x= -[log(12)/3log(3)]

OpenStudy (anonymous):

in g^x+2=12 I subtract both sides of equation by 2 so g^x=10 Again, log on both sides xlog(g)=log(10) thus log(g)=log(10)/x here you substitute the value of x.

OpenStudy (anonymous):

and that gives you log(g)= -3log(10)log(3)/log(12) then you take the anti log on both sides.

OpenStudy (anonymous):

I believe that solves your query.